OpenStudy (anonymous):

12d2 + 4d – 1 A. (6d + 1)(2d + 1) B. (6d – 1)(2d – 1) C. (6d – 1)(2d + 1) D. (6d + 1)(2d – 1) I think it is D

jimthompson5910 (jim_thompson5910):

no it's not D

OpenStudy (anonymous):

Am I correct?

OpenStudy (anonymous):

is it C

jimthompson5910 (jim_thompson5910):

good you got it

OpenStudy (anonymous):

thanks can you help me with one more please?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

A machinist creates a washer by drilling a hole through the center of a circular piece of metal. If the piece of metal has a radius of x + 10 and the hole has a radius of x + 6, what is the area of the washer? A. B. C. D.

jimthompson5910 (jim_thompson5910):

Area of washer = Area of metal - Area of drilled out hole btw, your answer choices are blank

OpenStudy (anonymous):

i noticed

OpenStudy (golikrishna):

12d2+4d-1 12d2+6d-2d-1 6d(2d+1)-1(2d+1) (6d-1)(2d+1)

OpenStudy (anonymous):

x^2-8piex-64pie x^2+8piex-64pie 8piex-64pie 8piex+64pie

OpenStudy (anonymous):

these are the answers

jimthompson5910 (jim_thompson5910):

Use the idea that A = pi*r^2 and plug in r = x+10 to find the area of the metal disk (the largest disk) Then plug in r = x+6 to find the area of the drilled out hole. Afterwards, subtract the two.

OpenStudy (anonymous):

so my is answer is going to be B. x^2+8piex-64pie

jimthompson5910 (jim_thompson5910):

what is (x+10)^2?

OpenStudy (anonymous):

100

jimthompson5910 (jim_thompson5910):

(x+10)^2 (x+10)(x+10) x(x+10)+10(x+10) x^2+10x+10x+100 x^2+20x+100 So (x+10)^2 = x^2+20x+100

OpenStudy (anonymous):

is the answer B.

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

then the answer has to be C then

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

a

jimthompson5910 (jim_thompson5910):

still no